OpenStudy (anonymous):

Select the correct graph of -x - 2y = 6

OpenStudy (usukidoll):

I don't see any answer choices given, but you should change your equation into y = mx+b form where m is the slope and b is the y-intercept.

OpenStudy (amoodarya):

find two point of the line ,and draw then compare with your choices for example x=-2 -(-2)-2y=6 -2y=4 y=-2 (-2,-2) x=0 0-2y=6 y=-3 (0,-3)